Main Street Service bought four Michelin tires from a wholesaler for $343.00 and sold the tires at a markup of 30% of the selling price. (a) For how much were the tires sold? (b) What is the rate of markup based on cost?


1
Expert's answer
2021-01-11T15:22:03-0500

a)

"C=\$343.00, M=S\times\dfrac{30\%}{100\%}=0.3S"

"S=C+M"

Substitute


"S=343+0.3S"

"0.7S=343"

"S=\$490"

b)


"M=0.3S=0.3(\$490)=\$147"

"\dfrac{M}{C}\cdot100\%=\dfrac{147}{343}\cdot100\%=42.857\%"

The rate of markup based on cost is "42.857\%."